OpenStudy (anonymous):

If a student has three test marks of 73% , 82% and 65% and wants an average of 80% over four tests, what must be her mark on the next test be ?

OpenStudy (anonymous):

\[(73+82+65+x)\div4=\ge80\] So if you do the math, then x, when solved, would be 100. Therefore she must get 100% on her next test to get an average of 80% or above.

OpenStudy (lilg132):

80 * 4 = 320 so all test mark percentages added together should equal to 320 so right now we have 73 + 82 + 65 = 220 320 - 220 = 100 so she needs 100% to get 80% average overall remember to click good answer :)